# Soundscape Captioning using Sound Affective Quality Network and Large Language Model

### v2

@@ III-A1 Mel-based branch
where $F_{i}$ denotes the $i$-th convolution layer’s RFS relative to the input feature map, $k$ denotes the convolution kernel size, and $stride$ defaults to 1. $F_{0}$ is the RFS of each point of the input feature map, i.e., $F_{0}=1$. If there is no pooling operation, according to Eq. (1), in the first Conv2D block, the RFS of the first layer on the time axis is $F_{1}=9$, and that of the second layer is $F_{2}=17$. For the dilated convolution, the RFS is
@@
where $r$ is the dilation rate. For the second Conv2D block with dilation rate (2, 1), on the time axis, the third layer’s RFS is $F_{3}=33$, and the fourth layer’s RFS is $F_{4}=49$. Similarly, for the third Conv2D block with dilation rate (3, 1), the fifth and sixth layers’ RFS on the time axis are 73 and 98, respectively. With these convolution blocks without pooling, SoundAQnet requires the input acoustic features to be at least 98 frames long. With the frame hop of 10ms, the corresponding input clip length is at least 980ms. It is challenging to identify AS or AE from 1-second audio clips, even for humans, let alone the 8D AQs. Furthermore, if pooling is not used in these Conv2D blocks, it will increase the parameters and the computation load. After comprehensive trade-offs, we add pooling operations to these multiscale convolution blocks, resulting in a minimum input audio length of 2.80s.
@@ III-A2 Loudness-based branch
The loudness of audio events affects human perception of them. Many models for noise annoyance within one class of sounds (e.g., road traffic) even exclusively rely on loudness. The Mel spectrogram can estimate the corresponding loudness changes to some extent. However, as perceived loudness has been studied in psychoacoustics for decades, it is more effective to directly introduce loudness as a psychoacoustic feature in model training instead of letting the model estimate it implicitly. In SoundAQnet, the loudness per time frame is extracted after calibration with a reference signal according to the ISO 532-1 standard [22]. The input size of this branch is ($N$, 1), where $N$ is the number of frames, and the multiscale convolution kernels of the loudness-based sub-branches are [($3$, 1), ($5$, 1), ($7$, 1), ($9$, 1)]. The rest of the loudness branch is the same as the Mel branch.
III-A3 Graph-based multiscale embedding fusion
@@
To fuse representations from these sub-branches, we consider the representation embeddings as node features and build a fully connected soundscape-dependent multiscale sound-AQ representation graph. Here, our hypothesis is that since the model is trained with co-supervised labels of AS, AE, and AQ in the soundscape, the sound-AQ representation graph will automatically couple the acoustic environment and AQ while updating node features and learning relationships between nodes with different time granularities. That is, by updating the features of edges connecting nodes, the message about the differences between different timescale nodes is passed to each other through edges in the graph, thereby further aggregating and fusing information from different scales. Thus, it is crucial to simultaneously learn edge features in the sound-AQ representation graph during updating. In ordinary graph convolution, only the node features are updated. Thus, we use the gated graph convolutional network (GatedGCN) [33] in the graph-based multiscale embedding fusion layer, where node and edge features are updated simultaneously. Once the soundscape-dependent sound-AQ representation graphs containing $n$ nodes and $n\times n$ edges are obtained, we use one layer of GatedGCN to model these graphs. Note that the number of nodes here is $n=8$, and the dimension of each node embedding is 64.
III-A4 Co-embedding and separate embedding layers
@@
To absorb the information updated by the sound-AQ graph-based fusion while considering its original acoustic representations, we residually connect [34] the node embeddings output by the graph with their representations before fusion. We concatenate all embeddings and input them into the common embedding layer to learn the acoustic context- and AQ-related embeddings. This allows the classification and regression tasks to use all the information within the common embedding captured by the model. Next, separate embedding layers are used for the ASC and AEC tasks, and the human-perceived AQ regression tasks, to learn representations for each target individually.
III-A5 Final loss
@@
As suggested in [35], the two axes of the SCM in Fig. 1, ISO Pleasantness (ISOP) and ISO Eventfulness (ISOE), can be calculated as follows:
$ISOP=k^{-1}(\sqrt{2}r_{pl}-\sqrt{2}r_{an}+r_{ca}-r_{ch}+r_{vi}-r_{mo})\vspace{-0mm}$ (3)
$ISOE=k^{-1}(\sqrt{2}r_{ev}-\sqrt{2}r_{ue}-r_{ca}+r_{ch}+r_{vi}-r_{mo})\vspace{-1mm}$ (4)
@@
where $r_{\{pl,ev,ch,vi,ue,ca,an,mo\}}$in $\{1,2,3,4,5\}$ are human response values corresponding to 8D AQs: pleasant, eventful, chaotic, vibrant, uneventful, calm, annoying, and monotonous, respectively, and $k=8+\sqrt{32}$. ISOP and ISOE are related to AQs, so the model’s prediction for ISOP can imply the overall performance of human-perceived AQ predictions.
@@
The SoundAQnet classifies 2 objectives (AS, AE) and regresses 10 objectives (ISOP, ISOE, 8D AQs). For ASC tasks, cross entropy (CE) [28] is used as the loss function measuring the difference between the prediction $p_{s}$ and its label ${y}_{s}$, i.e. $\mathcal{L}_{\text{1}}=CE(p_{s},y_{s})$. For AEC tasks, binary cross entropy (BCE) is used as the loss function measuring the difference between the prediction $p_{e}$ and its label ${y}_{e}$, i.e. $\mathcal{L}_{\text{2}}=BCE(p_{e},y_{e})$. For AQ regression tasks, mean squared error (MSE) is the loss function. Specifically, $\mathcal{L}_{\text{3}}=MSE(p_{isop},ISOP)$ and $\mathcal{L}_{\text{4}}=MSE(p_{isoe},ISOE)$, where $p_{isop}$ and $p_{isoe}$ are predictions of ISOP and ISOE, respectively. Also, $\mathcal{L}_{\text{n}}=MSE(p_{aq},y_{aq}),n\in[5,12]$, where $p_{aq}$ and ${y}_{aq}$ are predictions and labels of each type of AQ in 8D AQs. There are 12 losses in SoundAQnet, and optimizing the multiple objectives with multiple losses is challenging.
The typical Pareto optimization [36] in multi-objective optimization is unsuitable for SoundAQnet. Because the quantification of AQ has a certain degree of ambiguity, assuming that 3$\pm$0.25$\approx$3 for $r_{pl}$, its prediction $\pm$0.1 has little impact on the final AQ output. Hence, compared to emotion-related AQs, SoundAQnet needs to perform better in ASC and AEC with explicit classification goals, i.e., SoundAQnet does not aim to achieve the Pareto optimality of all 12 objectives. In addition, human perception times for various scenes, events, and emotions may vary, which means different rates, and the classification and regression losses in the 12 losses are of different natures. Hence, GradNorm-like optimizations [37], which aim to learn multiple tasks at a similar rate from a gradient view, do not suit SoundAQnet. After considering the computational effort and training speed, the uncertainty weighting [38] is adopted to fuse the 12 losses, as follows:
$\mathcal{L}=\sum\nolimits_{i=1}^{2}(\frac{1}{{\sigma_{i}}^{2}}{\mathcal{L}}_{i}+\text{log}{\sigma_{i}})+\sum\nolimits_{j=3}^{12}(\frac{1}{{2\sigma_{j}}^{2}}{\mathcal{L}}_{j}+\text{log}{\sigma_{j}})\vspace{-1mm}$ (5)
@@
where the learnable noise parameter $\sigma$ denotes the task’s uncertainty, and log is the penalty term that prevents $\sigma$ from becoming too large. The larger the uncertainty $\sigma$, the smaller the contribution of the loss to the overall loss.
@@
III-B Language model: A general LLM
For SoundSCaper in Fig. 3, the language model aims to convert the discrete information of AS and AE and the emotion-relevant AQ into a textual description of the soundscape, with the help of the prior knowledge embedded in LLM.
III-B1 Fusion strategy
In the proposed SoundSCaper system, the LLM serves two purposes: (1) to present the knowledge extracted from the audio recordings using SoundAQnet in fluent lay language descriptions; (2) to introduce common knowledge on expected audio events (AEs) in a specific context, i.e., a specific acoustic scene (AS). One could train a LLM on a large body of soundscape literature to achieve both purposes, but in the SoundSCaper system, a generic pre-trained GPT is used instead. This choice has several advantages: it avoids the time-consuming and laborious training of a soundscape LLM; directly using generic LLMs allows the acoustic and language models to be updated separately; the system can be extended with non-acoustic contextual information, if available; and captions can be easily provided in another language.
@@ III-B1 Fusion strategy
Strategies for integrating acoustic information into LLM can be divided into early fusion, mid fusion, and late fusion [39], corresponding to fuse at the feature level, intermediate representation, and decision level, respectively. In the pipeline in Fig. 5, according to $\boldsymbol{D}=lm(AS,AE,AQ)$ in Section II, the decision results of scenes, events, and affective qualities parsed by SoundAQnet are fed into GPT. The benefit of such decision fusion is that the proposed SoundSCaper pipeline can directly use general LLMs such as GPT.
III-B2 Related LLMs
According to OpenAI [18] services, this paper has three alternative LLMs: DaVinci, GPT-3.5-Turbo, and GPT-4. The soundscape captioning task mainly involves generating text descriptions based on the input. Hence, we choose GPT-3.5-Turbo, which offers a tradeoff among generation accuracy, response speed, number of tokens, and cost.
III-B3 Customized LLM for the task
As shown in Fig. 5, we integrate prompt engineering to enhance the output’s contextual accuracy, affective depth, and narrative clarity, to generate text descriptions for soundscape captioning tasks.
@@ III-B3 Customized LLM for the task
Token management: To optimize and control the input and output of LLM, the tokens are managed. For an audio clip, the scene, AS, is unique, but multiple AEs may be detected simultaneously. Hence, to process AE probabilities predicted by SoundAQnet, we first use an empirical threshold of 0.3 to obtain the text labels of AEs present in the audio clip. Then, for human-perceived AQs, we prioritize input tokens with strong responses, that is, high predicted values. This is because human attention is often attracted by the dominant AE while being influenced by the dominant AQ. The soundscape captioning task aims to describe the most relevant information in soundscapes. Additionally, we instruct the LLM to limit the generated descriptions to 200 tokens. These strategies manage the consumption of input and output tokens.
@@
Information priming: We offer the soundscape definition according to ISO 12913-1:2014 [1] to prime LLM with a conceptual framework to emphasize that perception (psychology) and understanding (cognition) should be included in the description, as well as context, people, and society.
@@
Chain-of-thought (CoT) prompting: This part guides the LLM through a logical analysis sequence, from AS and AE recognition to AQ evaluation. This structured approach aids in systematically tackling complex auditory and affective analyses. The task is decomposed into smaller, focused subtasks to help LLM understand the relationship between input acoustic environment information and AQs based on its large-scale prior knowledge to ensure the generation of comprehensive and accurate descriptions. The main prompts are as follows:
@@
Please see full prompts and scripts on the homepage.
IV Acoustic model experiment
IV-A Dataset
@@
ARAUS contains 25440 30-second binaural audio clips, totaling 212 hours. With the efforts of 605 experimental participants, each audio clip has 8D AQ values annotated according to ISO/TS 12913-2 [26]. ARAUS is augmented on the Urban Soundscapes of the World (USotW) [42] dataset. Each augmented soundscape is made by digitally adding maskers (birds, water, wind, traffic, construction, or silence) to an urban soundscape recording at soundscape-to-masker ratios [27]. The maskers are AEs. Hence, ARAUS meets the needs of SoundAQnet training with affective supervision information. Unfortunately, ARAUS does not have AS and AE labels.
@@
The USotW dataset, as ARAUS synthetic material, contains 360-degree video clips with precise locations, which allows us to easily identify their corresponding scene. There are 3 classes of ASs, namely {public square, park, street traffic}. Following ARAUS synthesis rules, we obtain the AS labels of ARAUS.
@@
Although six types of AEs have been explicitly added in ARAUS, we cannot directly use the six labels as AE labels because USotW already contains numerous AEs. To obtain the detailed AE labels in ARAUS, we first use the excellent audio pre-trained model PANNs [21] to label each audio clip with a one-second-level pseudo-label. Since the PANNs model is trained on AudioSet, a large-scale dataset with 527 classes of AEs, each one-second audio clip is assigned with a 527-dimensional soft pseudo label, corresponding to the probability of 527 classes of AEs within this second. Then, the soft pseudo-labels are mapped into hard pseudo-labels consisting of {0, 1} by a threshold of 0.5. After accumulating and sorting the hard pseudo-labels for all one-second segments, we obtain the number of occurrences for the 527 classes of AEs in ARAUS, ranked from high to low. After considering the six types of AEs added in ARAUS, a total of 15 AE labels are obtained, which are {Bird, Animal, Wind, Water, Natural sounds, Vehicle, Traffic, Sounds of things, Environment and background, Outside, rural or natural, Speech, Human sounds, Music, Noise, Silence}. However, for training SoundAQnet, only clip-level AE labels are needed to distinguish whether the target AE is within the input clip. Hence, we again use PANNs to label each audio clip’s clip-level 527 AE probabilities. Then, the probabilities of 15 classes of target AEs are taken out and binarized into hard labels using a threshold of 0.1.
@@
In the ARAUS experiment [27], the validation set has 5040 samples, while the test set has only 48 samples. The size of the test set may be too small to effectively evaluate the model performance. Thus, we randomly shuffled and re-divided the ARAUS dataset. In proportion, 19152 30-second binaural audio clips are randomly selected as the training set; 2520 and 3576 audio clips are chosen as the validation and test sets, respectively. To avoid the intersection between the three sets, the number of used audio clips is 25248, not 25440.
IV-B Experimental setup of acoustic model

### v3

@@ IV-A1 Network
where $F_{i}$ denotes the $i$-th convolution layer’s RFS relative to the input feature map, $F_{0}=1$, $k$ is the convolution kernel size, and stridestride defaults to 1. If there is no pooling operation, according to Eq. (1), in the first Conv2D block, the RFS of the first layer on the time axis is $F_{1}=9$, and that of the second layer is $F_{2}=17$. For the dilated convolution, the RFS is
@@
where $r$ is the dilation rate. For the second Conv2D block with dilation rate (2, 1), on the time axis, $F_{3}=33$, and $F_{4}=49$. For the third Conv2D block with dilation rate (3, 1), on the time axis, $F_{5}=73$, and $F_{6}=98$. With these Conv2D blocks without pooling, SoundAQnet requires the input features to be at least 98 frames long. With the frame hop of 10ms, the corresponding input clip length is at least 980ms. It is challenging to identify AS or AE from 1-second audio clips, even for humans, let alone the 8D AQs. Furthermore, if pooling is not used in Conv2D, it will increase the parameters and the computation load. After comprehensive trade-offs, we add pooling operations to these multiscale Conv2D blocks, resulting in a minimum input audio length of 2.80s.
@@
1.2) Loudness-based Branch: The loudness of AEs affects how humans perceive them, thereby affecting human-perceived AQ. Many models for noise annoyance within one class of sounds (e.g., road traffic) even rely exclusively on loudness. The Mel spectrogram can estimate the loudness changes to some extent. However, since perceived loudness has been studied in psychoacoustics for decades, it is more effective to directly use loudness as a psychoacoustic feature in the model instead of estimating it implicitly with the model. Loudness is introduced as a 1D feature with the unit of sone. In SoundAQnet, loudness features are extracted after calibration with a reference signal, i.e. a sine wave of 1kHz at 60dB, from the ISO 532-1 standard [20]. Given $N$ frames, the size of loudness features is ($N$, 1). Since the generation, development, and fading of emotions is a dynamic process with different time scales of importance, we also use multiscale convolution blocks in the loudness branch to extract the PAQs defined in ISO/TS 12913-3:2019 [4] shown in Fig. 1. More specifically, the multiscale convolution kernels used in the sub-branch are of dimension [($3$, 1), ($5$, 1), ($7$, 1), ($9$, 1)]. The remaining part of the loudness branch is the same as that in the Mel branch.
@@
1.3) Graph-based Multiscale Embedding Fusion: To fuse representations from Mel- and Loudness-based branches, we consider the representation embeddings as node features and build a fully connected soundscape-dependent multiscale sound-AQ representation graph. Here, our hypothesis is that since the model is trained with co-supervised labels of AS, AE, and AQ in the soundscape, the sound-AQ representation graph will automatically couple the acoustic environment and AQ while updating node features and learning relationships between nodes with different time granularities. That is, by updating the features of edges connecting nodes, the message about the differences between different timescale nodes is passed to each other through edges in the graph, thereby further aggregating and fusing information from different scales. Thus, it is crucial to learn edge features in the sound-AQ representation graph during updating. Then, we use the gated graph convolutional network (GatedGCN) [39] in the graph-based multiscale embedding fusion layer, where the node and edge features are updated simultaneously. Once the soundscape-dependent sound-AQ representation graphs containing $n$ nodes and $n\times n$ edges are obtained, we use one layer of GatedGCN to model these graphs. The number of nodes is $n=8$; the size of each node embedding is 64.
@@
1.4) Co-embedding and Separate Embedding Layers: To absorb the information updated by the sound-AQ graph while considering its original acoustic representations, we residually connect [40] the node embeddings output by the graph with the input to the graph. We concatenate all embeddings and input them into the common embedding layer to learn the acoustic context- and AQ-related embeddings. This allows the classification and regression tasks to use all the information within the common embedding captured by the model. Next, separate embedding layers are used for the ASC and AEC tasks, and the human-perceived AQ regression tasks, to learn representations for each target individually.
IV-A2 Loss Functions
@@
The SoundAQnet involves 2 classification objectives (AS, AE) and 10 regression objectives (ISOP, ISOE, 8D AQs). ISO Pleasantness (ISOP) and ISO Eventfulness (ISOE) in Fig. 1, can be calculated as follows:
$ISOP=k^{-1}(\sqrt{2}r_{pl}-\sqrt{2}r_{an}+r_{ca}-r_{ch}+r_{vi}-r_{mo})\vskip 0.0pt$ (3)
$ISOE=k^{-1}(\sqrt{2}r_{ev}-\sqrt{2}r_{ue}-r_{ca}+r_{ch}+r_{vi}-r_{mo})\vskip-2.84526pt$ (4)
@@
where $r_{\{pl,ev,ch,vi,ue,ca,an,mo\}}$ $\in$ $\{1,2,3,4,5\}$ are human response values to 8D AQs: pleasant, eventful, chaotic, vibrant, uneventful, calm, annoying, and monotonous, respectively, and $k=8+\sqrt{32}$. ISOP and ISOE are related to AQs, so the model’s prediction for ISOP can imply the overall performance of human-perceived AQ predictions.
2.1) AS: For the ASC tasks, cross entropy (CE) [15] is used as the loss function that measures the difference between the prediction $p_{s}$ and its label ${y}_{s}$, i.e. $\mathcal{L}_{\text{1}}=CE(p_{s},y_{s})$.
2.2) AE: For the AEC tasks, binary cross entropy (BCE) is used as the loss function that measures the difference between the prediction $p_{e}$ and its label ${y}_{e}$, i.e. $\mathcal{L}_{\text{2}}=BCE(p_{e},y_{e})$.
@@
2.3) AQ: For the AQ regression tasks, mean squared error (MSE) is used as the loss function. Specifically, $\mathcal{L}_{\text{3}}=MSE(p_{isop},ISOP)$ and $\mathcal{L}_{\text{4}}=MSE(p_{isoe},ISOE)$, where $p_{isop}$ and $p_{isoe}$ are predictions of ISOP and ISOE, respectively. Also, $\mathcal{L}_{\text{n}}=MSE(p_{aq},y_{aq}),n\in[5,12]$, where $p_{aq}$ and ${y}_{aq}$ are predictions and labels of each type of AQ in 8D AQs.
[…]
$\mathcal{L}=\sum\nolimits_{i=1}^{2}(\frac{1}{{\sigma_{i}}^{2}}{\mathcal{L}}_{i}+\text{log}{\sigma_{i}})+\sum\nolimits_{j=3}^{12}(\frac{1}{{2\sigma_{j}}^{2}}{\mathcal{L}}_{j}+\text{log}{\sigma_{j}})\vskip-2.84526pt$ (5)
@@
where the learnable noise parameter $\sigma$ denotes the uncertainty associated with the corresponding task [43], and the logarithmic function based penalty term prevents $\sigma$ from becoming excessively high. The higher the uncertainty $\sigma$, the lower the contribution of the loss to the total loss.
IV-B Language Model: Customised LLM for Caption Generation
This section explains how the AS, AE, and AQ information, obtained by SoundAQnet, can be turned into captions to describe the soundscape in human-understandable language by customizing an LLM with prompt engineering techniques.
@@ IV-B Language Model: Customised LLM for Caption Generation
IV-B1 Pipeline for Caption Generation
@@
As shown in Fig. 5, we design a pipeline to convert the numerical information of AS and AE and the emotion-relevant AQ into a textual description of the soundscape, with the help of the prior knowledge learned by an LLM. There are various methods for integrating acoustic information into LLM, including early, mid, and late fusion [44], corresponding to information fusion at the feature, intermediate representation, and decision level, respectively. In this work, we choose GPT as the LLM to fuse the decision results of scenes, events, and affective qualities parsed by SoundAQnet, i.e. $\boldsymbol{D}=lm(AS,AE,AQ)$ as discussed in Section III, where $lm$ represents the GPT model, such as DaVinci, GPT-3.5-Turbo, and GPT-4, according to OpenAI [18] services. We choose GPT-3.5-Turbo, which offers a tradeoff among generation accuracy, response speed, number of tokens, and cost. An advantage of this pipeline is that SoundSCaper’s pipeline can directly use general LLMs such as GPT and potentially benefit from their future updates.
In the SoundSCaper system, LLM serves two purposes: (1) to present knowledge extracted from audio recordings using SoundAQnet in fluent language; (2) to introduce common knowledge on expected AEs and AQs in a specific context, that is, a specific AS. Currently, there is no large-scale dataset in the soundscape domain that pairs audio with affective descriptive captions encompassing AE, AS, and human-perceived AQ. As a result, existing resources are inadequate for training audio-to-text soundscape-language models or for fine-tuning LLMs. In addition, fine-tuning LLMs on expert-annotated affective captions may introduce biases related to individual personalities and cultural perspectives. Moreover, fine-tuning LLM may increase the chance of hallucination [45]. For these reasons, SoundSCaper uses a general pre-trained GPT instead of fine-tuning a model. This choice has several advantages: it avoids the dilemma of no paired data of sound and affective description to support the training or fine-tuning of soundscape LLM, and also avoids time-consuming and laborious training; directly using generic LLMs allows the acoustic and language models to be updated separately; the system can be extended with non-acoustic contextual information, if available; and captions can be easily provided in other languages.
IV-B2 Customized LLM
As a novel design for soundscape caption generation, we integrate prompt engineering to enhance the output’s contextual accuracy, affective depth, and narrative clarity to decrease hallucinations, as shown in Fig. 5. The employed prompt engineering strategies include structured output generation, system prompt enhancement, semantic layer integration, and chain-of-thought (CoT) prompting, which are more effective than fine-tuning LLM in improving model performance and reducing hallucinations [46].
@@ IV-B2 Customized LLM
2.1) Token Management: To optimize the amount of input and output of the LLM, tokens are managed. For an audio clip, the scene, AS, is unique, but multiple AEs may be detected simultaneously. Hence, to process AE probabilities predicted by SoundAQnet, we first use an empirical probability threshold of 0.3 to obtain the text labels of AEs present in the audio clip. Then, for AQs predicted by SoundAQnet, we prioritize input tokens with strong responses, that is, high predicted values. This is because human attention is often attracted by the dominant AE while being influenced by the dominant AQ. The ASSC task aims to describe the most relevant information in soundscapes. Additionally, we instruct the LLM to limit the generated descriptions to 200 tokens. These strategies manage the consumption of input and output tokens.
@@
2.2) System Prompt Enhancement: We use the soundscape definition from ISO 12913-1:2014 [3] to provide LLM with a conceptual framework that incorporates perception (psychology) and understanding (cognition) in the description, as well as context, people, and society. In addition, the prompt asks the LLM to define its role as a soundscape expert to fit the soundscape research field.
2.3) Semantic Layer Integration: Rule-based semantic constraints are used to reduce hallucinations [46] in ASSC. The LLM is instructed to use only the provided AE labels and avoid interpretive or embellished descriptions. For example, vehicle should not be expressed as ‘the hum of engines and honking of horns’; LLM should directly list the AEs. These constraints help ground the output, improving accuracy and consistency without requiring LLM fine-tuning.
@@
2.4) Chain-of-thought Prompting: This part guides the LLM through a logical analysis sequence for structured output generation, from AS and AE classification to AQ regression. The structured approach aids in systematically tackling complex auditory and affective analyses. The task is decomposed into focused subtasks to help LLM understand the relationship between input acoustic environment information and AQs based on its large-scale prior knowledge to ensure comprehensive and accurate caption generation. Prompts are as follows:
@@
Full prompts and scripts can be found in the homepage.
V Acoustic model (SoundAQnet) experiment
V-A Dataset
@@
ARAUS contains 25440 30-second binaural audio clips, totaling 212 hours. With the efforts of 605 participants, each audio clip has 8D AQ values annotated according to ISO/TS 12913-2 [35]. ARAUS is augmented on the Urban Soundscapes of the World (USotW) [49] dataset. Each augmented soundscape is made by digitally adding maskers (birds, water, wind, traffic, construction, or silence) to an urban soundscape recording at soundscape-to-masker ratios [32]. The maskers are AEs. Hence, ARAUS meets the needs of SoundAQnet training with affective supervision information. Unfortunately, ARAUS does not have AS and AE labels.
@@
The USotW [49], as synthesis material for ARAUS, contains 360-degree video clips with GPS locations, which allows us to easily identify their corresponding scene. There are 3 classes of ASs, namely {public square, park, street traffic}. Following ARAUS synthesis rules, we obtain the AS labels of ARAUS.
@@
Although six types of AEs have been explicitly added in ARAUS, we cannot directly use the six labels as AE labels because USotW already contains numerous AEs. To obtain the detailed AE labels in ARAUS, we first use the excellent pre-trained audio model PANNs [12] to label each audio clip with a one-second-level pseudo-label. Since the PANNs model is trained on AudioSet, a large-scale dataset with 527 classes of AEs, each one-second audio clip is assigned with a 527-dimensional soft pseudo label, corresponding to the probability of 527 classes of AEs within this second. Then, the soft pseudo-labels are mapped into hard pseudo-labels consisting of {0, 1} by a threshold of 0.5. After accumulating and sorting the hard pseudo-labels for all one-second segments, we obtain the number of occurrences for the 527 classes of AEs in ARAUS, ranked from high to low. After considering the six types of AEs added in ARAUS, a total of 15 AE labels are obtained, which are {Bird, Animal, Wind, Water, Natural sounds, Vehicle, Traffic, Sounds of things, Environment and background, Outside, rural or natural, Speech, Human sounds, Music, Noise, Silence}. For training SoundAQnet, only clip-level AE labels are needed to distinguish whether the target AE is within the input clip. Hence, we again use PANNs to label the clip-level 527 AE probabilities for each audio clip. Then, the probabilities of 15 classes of target AEs are taken out and binarized into hard labels using a threshold of 0.1.
@@
In the experiment of ARAUS [32], the validation set has 5040 samples, while the test set has only 48 samples. The test set may be too small to effectively evaluate the performance of our model. Thus, we randomly shuffled and re-divided the ARAUS dataset. In proportion, 19152 30-second audio clips are randomly selected as the training set; 2520 and 3576 audio clips are chosen as the validation and test sets, respectively. To avoid intersections between the three sets, the number of audio clips used in this paper is 25248, rather than 25440.
V-B Experimental Setup of Acoustic Model